OpenStudy (anonymous):

find the slope (3,-20) (5,8)

OpenStudy (anonymous):

Do you know how to find the slope of a line?

OpenStudy (anonymous):

yes not very well,

OpenStudy (anonymous):

Okay do you know the formula?

OpenStudy (anonymous):

y1 y2 /x1 x2

OpenStudy (anonymous):

its not like you have written its m=(y2-y1)/(x2-x1) can you solve it now?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I got it but my teacher was like it doesn't matter which one goes first either one is fine she say a lot of ppl say that's how its pose to be .. but yeah ima try it

OpenStudy (anonymous):

uhh I got-160 over 24 shouldn't I do something else

OpenStudy (anonymous):

how you got it ? its wrong you have to subtract not to multiply

OpenStudy (anonymous):

oh I saw the x ... lol hold on

OpenStudy (anonymous):

What your teacher means is that it does matter which one of the x values is x1 and x2 all that matters is which ever value you use you us on the bottom you use on the top (or the y) so x1 3 or 5 but if you pick 3 for x1 you have to pick -20 for y1.

OpenStudy (anonymous):

-28 over 25

OpenStudy (anonymous):

5

OpenStudy (anonymous):

no no -2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so -28 over -2

OpenStudy (anonymous):

that can't be the complete answer can it?

OpenStudy (anonymous):

divide the two numbers

OpenStudy (anonymous):

14 :D ?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

I just found this website just now I thought it wouldn't answer my question until I do a bunch of stuff and popups I actually really liking this website im enjoying learning a lot I have a math teacher that pauses in the middle of sentences out of no where it's hard to learn my teachers sweet but be off subject alot

OpenStudy (anonymous):

Welcome to the site then. When I first found this site I was enjoying it just as much.

OpenStudy (anonymous):

:) I made a friend

OpenStudy (anonymous):

wanna help me find x in 12+5 -8= 12 x -10

OpenStudy (anonymous):

sure send me a message about it then you can close this question.

OpenStudy (anonymous):

12+5x-8=12x-10